# Quantum dust collapse with cosmological constant and methods for constructing conformal diagrams

### v2

@@ II Dust Loop Quantum Cosmology with Cosmological Constant
While the system under consideration has already been described on the genuine quantum level Husain and Pawlowski (2011), its particular application here (gluing to classical exterior metric) requires us to reduce our description to just quantum trajectories (of the expectation values). Furthermore, by its very nature, the consistency of the gluing requires us to restrict our studies to very sharply localized semiclassical states. Therefore, for the purpose of this work, it is sufficient to employ the so-called effective dynamics Singh and Vandersloot (2005) (which in turn is the 00th order cutoff of the more systematic semiclassical treatment Bojowald and Skirzewski (2006)). In this we follow exactly the treatment of Bobula and Pawłowski (2023). First, we approximate classical treatment by replacing the component operators with their expectation values, assuming further that the (expectation values of the) holonomy components can be represented as exponentiated momentum $\dutchcal{b}$, which gives $\dutchcal{v}:=\langle\hat{\dutchcal{v}}\rangle$ and $\sin\dutchcal{b}=\langle(i/2)(\hat{\mathcal{N}}^{2}-\hat{\mathcal{N}}^{-2})\rangle$, $\cos\dutchcal{b}=\langle(1/2)(\hat{\mathcal{N}}^{2}+\hat{\mathcal{N}}^{-2})\rangle)$. The chosen variables form a Poisson algebra with the bracket $\{v,b\}=-2/\hbar$ and the effective Hamiltonian, built by replacing all the component operators with listed variables/functions of $\dutchcal{b}$, takes the form
@@ III Exterior metric
where $M_{\mathrm{eff}}:=8\alpha_{o}r_{b}^{3}\dutchcal{v}_{\mathrm{in}}(A-C)(A-2C)/(9G)$. Throughout the rest of our work, we will restrict our analysis to the situation when (18) has 3 positive roots, say $X_{1}<X_{2}<X_{3}$ — this corresponds to the situation when $M_{\mathrm{eff}}\gg m_{\mathrm{Pl}}$ is comparable to the mass of collapsing matter in astrophysical scenarios and $\Lambda$ is close to observational values — see Appendix A for the complete discussion. In other words, we consider a dust ball collapse “slightly perturbed” [^4] This is not the perturbation in the sense that the system is linearized. We simply look for the dynamics when $\Lambda$ is small. by the presence of the cosmological constant. For notational purposes we also define $T_{1}:=T(X_{1})$, $T_{2}:=T(X_{2})$, and $T_{3}:=T(X_{3})$ according to (17). The exterior geometry is completely determined now. The line element (8) (with $F(X)$ given in (18)) suggest that the exterior describes black-hole-like, asymptotically de Sitter spacetime. We will indeed confirm that statement in the next sections.
@@ IV Methods for constructing conformal diagrams
  • Let $\dutchcal{s}$ be a monotonic and compatible with spherical symmetetry function at $\Omega$. In other words, $\Omega$ can be seen as a “tube” where values of $\dutchcal{s}$ label 2-spheres. For example, the “parameter” $T$ from Section III is such a function. Based on junction conditions, solve for the “trajectory” of $\Omega$ with respect to both coordinate systems, namely, calculate $(u_{*}^{\mathrm{in}}(\dutchcal{s}),v_{*}^{\mathrm{in}}(\dutchcal{s}))$ and $(u_{*}^{\mathrm{out}}(\dutchcal{s}),v_{*}^{\mathrm{out}}(\dutchcal{s}))$. Find inverses, that is $\dutchcal{s}(u_{*}^{\mathrm{out}})$ and $\dutchcal{s}(v_{*}^{\mathrm{out}})$. Note that the functions may be only piecewise invertible, however, the construction will still work. These coordinates will provide an abstract chart on the to-be Penrose-Cartan diagram through an assumption of them being dragged (constant) along null geodesics, forming the lines at $45^{o}$ angle on the diagram.
@@
    We aim to analyse the behaviour of an affine parameter at $u_{*}^{\mathrm{out}}=const.$ and $v_{*}^{\mathrm{out}}=const.$ families of null geodesics. We stress that the points where the affine parameter reaches infinite value indicate the spacetime endpoints. However, the spacetime may end in a finite value of the affine parameter while traversing along null geodesic — for example, a light ray can hit a singularity. Consider the following. Suppose we analyze $v_{*}^{\mathrm{out}}=const.$ geodesic, and let $u_{*}^{\mathrm{out}}$ be a parameter. The tangent vector to that geodesic, say $k^{\alpha}$, satisfies $k^{\alpha}_{;\beta}k^{\alpha}=\frac{\mathrm{d}\mathcal{F}^{out}(u_{*}^{\mathrm{out}})}{\mathrm{d}u_{*}^{\mathrm{out}}}\frac{1}{\mathcal{F}^{\mathrm{out}}(u_{*}^{\mathrm{out}})}k^{\alpha}$. It tells us that $u^{\mathrm{out}}$ is not affine, however, it is related to the one, say $\lambda$, via (see Chapter 1.3 of Poisson (2009) for the general discussion)
@@ V.2 Extension of the exterior coordinates
where $h$ is so-called pre-squishing function (see Section 5.2 of Schindler and Aguirre (2018) for explicit form and more details), $\epsilon_{u^{\text{out}}}$ and $\epsilon_{v^{\text{out}}}$ take values in $\{-1,1\}$ determining orientation of a block (a basic building component [^7] For example a conformal diagram of Schwarzschild spacetime consists of 4 blocks: black hole interior, white hole interior, a universe and a parallel universe. of a conformal diagram, see Section 4.4 of Schindler and Aguirre (2018) for a rigorous definition), constants $\tilde{c}_{u^{\text{out}}}$ and $\tilde{c}_{v^{\text{out}}}$ are responsible the location of each block. That is, in each block, constants present in (32) are not the same. Having that setup, it is straightforward to follow the rules of Schindler and Aguirre (2018) to plot the exterior parts (of our interest) of the conformal diagram. However, since we are interested in covering also the interior with (32) we need to extend the coordinates. We note that the parametrized trajectory of the surface of the collapsing dust ball in new coordinates (32) is the pair

### v3

@@ II Dust Loop Quantum Cosmology with Cosmological Constant
While the system under consideration has already been described on the genuine quantum level Husain and Pawlowski (2011), its particular application here (gluing to classical exterior metric) requires us to reduce our description to just quantum trajectories (of the expectation values). Furthermore, by its very nature, the consistency of the gluing requires us to restrict our studies to very sharply localized semiclassical states. Therefore, for the purpose of this work, it is sufficient to employ the so-called effective dynamics Singh and Vandersloot (2005) (which in turn is the 0th order cutoff of the more systematic semiclassical treatment Bojowald and Skirzewski (2006)). In this we follow exactly the treatment of Bobula and Pawłowski (2023). First, we approximate classical treatment by replacing the component operators with their expectation values, assuming further that the (expectation values of the) holonomy components can be represented as exponentiated momentum $\dutchcal{b}$, which gives $\dutchcal{v}:=\langle\hat{\dutchcal{v}}\rangle$ and $\sin\dutchcal{b}=\langle(i/2)(\hat{\mathcal{N}}^{2}-\hat{\mathcal{N}}^{-2})\rangle$, $\cos\dutchcal{b}=\langle(1/2)(\hat{\mathcal{N}}^{2}+\hat{\mathcal{N}}^{-2})\rangle)$. The chosen variables form a Poisson algebra with the bracket $\{v,b\}=-2/\hbar$ and the effective Hamiltonian, built by replacing all the component operators with listed variables/functions of $\dutchcal{b}$, takes the form
@@ III Exterior metric
where $M_{\mathrm{eff}}:=8\alpha_{o}r_{b}^{3}\dutchcal{v}_{\mathrm{in}}(A-C)(A-2C)/(9G)$. As an aside, the metric (18) was also found in Lin and Zhang (2024) by solving loop quantum effective equations of motion in spherical symmetry accounting for the presence of a cosmological constant in midisuperspace approach. This result further supports the viability of the junction conditions formalism we employed. Throughout the rest of our work, we will restrict our analysis to the situation when (18) has 3 positive roots, say $X_{1}<X_{2}<X_{3}$ — this corresponds to the situation when $M_{\mathrm{eff}}\gg m_{\mathrm{Pl}}$ is comparable to the mass of collapsing matter in astrophysical scenarios and $\Lambda$ is close to observational values — see Appendix A for the complete discussion. In other words, we consider a dust ball collapse “slightly perturbed” [^4] This is not the perturbation in the sense that the system is linearized. We simply look for the dynamics when $\Lambda$ is small. by the presence of the cosmological constant. For notational purposes we also define $T_{1}:=T(X_{1})$, $T_{2}:=T(X_{2})$, and $T_{3}:=T(X_{3})$ according to (17). The exterior geometry is completely determined now. The line element (8) (with $F(X)$ given in (18)) suggest that the exterior describes black-hole-like, asymptotically de Sitter spacetime. We will indeed confirm that statement in the next sections.
@@ IV Methods for constructing conformal diagrams
  • Let $\dutchcal{s}$ be a monotonic and compatible with spherical symmetry function at $\Omega$. In other words, $\Omega$ can be seen as a “tube” where values of $\dutchcal{s}$ label 2-spheres. For example, the “parameter” $T$ from Section III is such a function. Based on junction conditions, solve for the “trajectory” of $\Omega$ with respect to both coordinate systems, namely, calculate $(u_{*}^{\mathrm{in}}(\dutchcal{s}),v_{*}^{\mathrm{in}}(\dutchcal{s}))$ and $(u_{*}^{\mathrm{out}}(\dutchcal{s}),v_{*}^{\mathrm{out}}(\dutchcal{s}))$. Find inverses, that is $\dutchcal{s}(u_{*}^{\mathrm{out}})$ and $\dutchcal{s}(v_{*}^{\mathrm{out}})$. Note that the functions may be only piecewise invertible, however, the construction will still work. These coordinates will provide an abstract chart on the to-be Penrose-Cartan diagram through an assumption of them being dragged (constant) along null geodesics, forming the lines at $45^{o}$ angle on the diagram.
@@
    We aim to analyse the behaviour of an affine parameter at $u_{*}^{\mathrm{out}}=const.$ and $v_{*}^{\mathrm{out}}=const.$ families of null geodesics. We stress that the points where the affine parameter reaches infinite value indicate the spacetime endpoints. However, the spacetime may end in a finite value of the affine parameter while traversing along null geodesic — for example, a light ray can hit a singularity. Consider the following. Suppose we analyze $v_{*}^{\mathrm{out}}=const.$ geodesic, and let $u_{*}^{\mathrm{out}}$ be a parameter. The tangent vector to that geodesic, say $k^{\alpha}$, satisfies $k^{\alpha}_{;\beta}k^{\beta}=\frac{\mathrm{d}\mathcal{F}^{out}(u_{*}^{\mathrm{out}})}{\mathrm{d}u_{*}^{\mathrm{out}}}\frac{1}{\mathcal{F}^{\mathrm{out}}(u_{*}^{\mathrm{out}})}k^{\alpha}$. It tells us that $u^{\mathrm{out}}$ is not affine, however, it is related to the one, say $\lambda$, via (see Chapter 1.3 of Poisson (2009) for the general discussion)
@@
    Note that the extension through the relation $u^{\rm in}_{*}\leftrightarrow\dutchcal{s}\leftrightarrow u^{\rm out}_{*}$ and $v^{\rm in}_{*}\leftrightarrow\dutchcal{s}\leftrightarrow v^{\rm out}_{*}$ is realized for each coordinate separately. As illustrated in Figure 1, given a spacetime point of our interest, namely $p_{\dutchcal{s}_{12}}$ coordinatized by the pair $(u_{*}^{\mathrm{out}}\big{|}_{p_{\dutchcal{s}_{12}}},v_{*}^{\mathrm{out}}\big{|}_{p_{\dutchcal{s}_{12}}})$ and located in the second componential spacetime, there is only one, say ”right-moving”, $u_{*}^{\mathrm{out}}=const.$ null geodesic connecting $p_{\dutchcal{s}_{12}}$ with $p_{\dutchcal{s}_{2}}$ at $\Omega$. Therefore, we have unique value $\dutchcal{s}=\dutchcal{s}_{2}$ (and then $u_{*}^{\mathrm{in}}(\dutchcal{s}=\dutchcal{s}_{2}))$ corresponding to that $u_{*}^{\mathrm{out}}\big{|}_{p_{\dutchcal{s}_{12}}}=u_{*}^{\mathrm{out}}\big{|}_{p_{\dutchcal{s}_{2}}}=const.$ null geodesic. Similarly there is only one, say ”left-moving”, $v_{*}^{\mathrm{out}}=const.$ null geodesic connecting $p_{\dutchcal{s}_{12}}$ with $p_{\dutchcal{s}_{1}}$ at $\Omega$. Therefore there is only one $\dutchcal{s}=\dutchcal{s}_{1}$ (and then $v_{*}^{\mathrm{in}}(\dutchcal{s}=\dutchcal{s}_{1}))$) corresponding to that $v_{*}^{\mathrm{out}}\big{|}_{p_{\dutchcal{s}_{12}}}=v_{*}^{\mathrm{out}}\big{|}_{p_{\dutchcal{s}_{1}}}=const.$ null geodesic.
@@ V.2 Extension of the exterior coordinates
where $h$ is so-called pre-squishing function (see Section 5.2 of Schindler and Aguirre (2018) for explicit form and more details), $\epsilon_{u^{\text{out}}}$ and $\epsilon_{v^{\text{out}}}$ take values in {$,1\}$ 1 } determining orientation of a block (a basic building component [^7] For example a conformal diagram of Schwarzschild spacetime consists of 4 blocks: black hole interior, white hole interior, a universe and a parallel universe. of a conformal diagram, see Section 4.4 of Schindler and Aguirre (2018) for a rigorous definition), constants $\tilde{c}_{u^{\text{out}}}$ and $\tilde{c}_{v^{\text{out}}}$ are responsible the location of each block. That is, in each block, constants present in (32) are not the same. Having that setup, it is straightforward to follow the rules of Schindler and Aguirre (2018) to plot the exterior parts (of our interest) of the conformal diagram. However, since we are interested in covering also the interior with (32) we need to extend the coordinates. We note that the parametrized trajectory of the surface of the collapsing dust ball in new coordinates (32) is the pair